x^2 + 3x - 10 >= 0
Answer by roseo(33)

Did you want the solutions for x^2 + 3x - 10>0 if so then solve the quadratic x^2+3x-10=0 by factoring.
(x+5)(x-2) = 0
x+5=0 and x-2=0 so
x=-5 and x=2 so we must now examine the intervals (negative infinity, -5), (-5,2)and (2, infinity). To do so pick a value in each interval and see which ones satisfy the inequality.
1st interval pick -10 so (-10)^2+3(-10)-10 = 60 which is > zero so this interval works
2nd interval pick 0 - o^2 +3(0)-10 = -10 which is not > zero so this interval does not work
3rd interval pick 3 - 3^2+3(3)-10 = 8 which is > zero so this interval works. So our solution set would contain (negative infinity, -5)and (2, infinity)